Walt made an extra $10,000 last year from a part-time job. He invested part of the money at 8% and the rest at 6%. He made a total of $720 in interest. How much was invested at 6%?
Answer by mananth(16949)

Investment II 8.00% per annum ---x
Investment II 7.00% per annum ---y

x+y= 10000 ------------------------1
8.00% x + 7.00% y = $720.00
Multiply by 100
8 x + 7 y = $72,000.00 --------2

Multiply (1) by -8
we get

-8x-8y =-80000.00

Add this to (2)

0 x -1 y = -$8,000.00

divide by -1

y = $8,000.00 investment at 7.00%
Balance $2,000.00 investment at 8.00%
CHECK
$2,000.00 --------- 8.00% ------- $160.00
$8,000.00 ------- 7.00% ------- $560.00
Total ------- $720.00